OpenStudy (anonymous):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (anonymous):

is it the a+ar+ar^2... formula?

OpenStudy (anonymous):

a+ar+ar^2 is short form for ar^0+ar^1+ar^2 .... ar^7 1 2 3 8th will be

OpenStudy (anonymous):

a is just 8 since it is first term a=8 r is ar^7=781250 -> 8*r^7=781250 solve for r

OpenStudy (anonymous):

r=5.162

OpenStudy (anonymous):

knowing that you then use this formula http://www.regentsprep.org/Regents/math/algtrig/ATP2/GeoSeq.htm to find the sum :D

OpenStudy (anonymous):

(8*(1-5.162^8)/(1-5.162)

OpenStudy (anonymous):

roughly 969016.23049467104062767206

OpenStudy (campbell_st):

just a slight correction to the above work \[a_{n} = a_{1} \times r^{n -1}\] you know a1 = 10 and an = 781250 and n = 8 so you have \[781250 = 10 \times r^{8 -1}\] divide both sides of the equation by 10 \[78125 = r^7\] to find the common ratio r \[\sqrt[7]{78125} = r\] so in actual fact r = 5 for the sum of a geometric series \[S_{8} = \frac{10(1 - 5^8)}{1 - 5}\] evaluate for the answer. Hope this helps.